# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 2 Background
Using birthpoints and deathpoints, the $L$-fibered barcode of an interval decomposable multi-parameter persistence module ${\mathbb{M}}=\bigoplus_{i\in\mathcal{I}}I_{i}$ is written as:
@@
${\mathcal{F}}{\mathcal{B}}({\mathbb{M}})_{L}=\{{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}):l\in L\}=\{([b^{I_{i}}_{l},d^{I_{i}}_{l}])_{i\in\mathcal{I}}:l\in L\}.$ (3)
3 Approximate candidate decompositions
In this section, we present our family of topological invariants for multi-parameter persistence modules, defined as candidate decompositions into interval summands, in Section 3.1. We then prove their approximation guarantees in Section 3.2.
3.1 Candidate modules and interval decompositions
@@
In this section, given a multi-parameter persistence module ${\mathbb{M}}$, we define a family of good proxies for ${\mathbb{M}}$ (see Definition 3.3), called candidate decompositions. These candidate decompositions are multi-parameter persistence modules that can be decomposed into intervals summands, and approximate ${\mathbb{M}}$ well when ${\mathbb{M}}$ is interval decomposable itself, in the sense that ${d_{\rm I}}({\mathbb{M}},\tilde{\mathbb{M}})$ and ${d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}})$ are upper bounded for any candidate decomposition $\tilde{\mathbb{M}}$. Practical computations of instances of such candidate decompositions can be done with our MMA algorithm, that we provide in Section 4.2.
(Exact) matchings and grids of lines.
@@
Our candidate decompositions depend on so-called matching functions and $\delta$-grids of lines, that we now define.
Definition 3.1 ((Exact) matching).
Let ${\mathbb{M}}$ be a multi-parameter persistence module. Let $l,l^{\prime}$ be two lines in $\mathbb{R}^{n}$. A map $\sigma$ between the persistence barcodes $\sigma\colon{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})\rightarrow{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}})\cup\left\{\varnothing\right\}$ is called an (${\mathbb{M}}$-)matching between $l$ and $l^{\prime}$ if the restriction of $\sigma$ to $\sigma^{-1}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}}))$ is injective. Furthermore, if ${\mathbb{M}}$ is interval decomposable, ${\mathbb{M}}=\bigoplus_{i\in\mathcal{I}}I_{i}$, we let $\iota_{1}:{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})\rightarrow\mathcal{I}$ and $\iota_{2}:{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}})\rightarrow\mathcal{I}$ be maps that retrieve the indices of the summands corresponding to the bars in ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}})$ and ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}})$ as per Equation (3). Then, $\sigma$ is said to be exact if $\iota_{1}(b)=\iota_{2}(\sigma(b))$ for any bar $b\in\sigma^{-1}({\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l^{\prime}}}))$. In other words, bars that are matched under $\sigma$ correspond to the same underlying interval summand of ${\mathbb{M}}$. Note that the interval decomposition of ${\mathbb{M}}$ always induces a canonical exact matching, that we denote $\sigma_{\mathbb{M}}$.
Definition 3.2 ($\delta$-grid of lines).
@@
where $l_{\delta\cdot u}:=\delta\cdot u+\boldsymbol{e}_{\Delta}$ is the diagonal line with direction vector $\boldsymbol{e}_{\Delta}=[1,\dots,1]^{T}\in\mathbb{R}^{n}$ passing through $\delta\cdot u$.
Candidates decompositions.
Our candidate decompositions are, roughly speaking, modules with the same fibered barcodes than ${\mathbb{M}}$ on a $\delta$-grid of lines. When ${\mathbb{M}}$ is interval decomposable, we also ask the exact matchings induced by the candidate decompositions to commute with the one induced by ${\mathbb{M}}$ (we discuss such matchings in Section 4.4).
Definition 3.3 (Candidate).
Let ${\mathbb{M}}$ be a multi-parameter persistence module. Let $K$ be a hyperrectangle in $\mathbb{R}^{n}$ containing the multigraded Betti numbers of ${\mathbb{M}}$  [^2] See [BL22, Section 4.2] for the definition of multigraded Betti numbers. Note that understanding multigraded Betti numbers is not required to follow this work. Moreover, in practice, when ${\mathbb{M}}$ is computed from the persistent homology of a multi-parameter filtration on a simplicial complex $S$ induced by a continuous function $f$, it suffices to find a hyperrectangle $K$ which contains all the multi-parameter filtration values, i.e., such that $\{f(\sigma):\sigma\in S\}\subseteq K$., and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}=\{x\in\mathbb{R}^{n}:d_{\infty}(x,K)\leq\delta\}$, where $d_{\infty}$ stands for the $\left\lVert\cdot\right\rVert_{\infty}$ distance. A multi-parameter persistence module $\tilde{\mathbb{M}}_{\delta}$ is called a $\delta$-candidate decomposition of ${\mathbb{M}}$ if
  1. $(i)$
    $\tilde{\mathbb{M}}_{\delta}$ is interval decomposable: $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}\tilde{I}_{i}$,
  2. $(ii)$
    ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})={\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big{|}}\right|_{l}})$ for any $l\in L$, i.e., their $L$-fibered barcodes are the same, and
  3. $(iii)$
    if ${\mathbb{M}}=\bigoplus_{i\in\mathcal{I}}I_{i}$ is interval decomposable, then there exists a bijection $\nu\colon\mathcal{I}_{L}\to\tilde{\mathcal{I}}$, where $\mathcal{I}_{L}=\{i\in\mathcal{I}:I_{i}\cap L\neq\varnothing\}$, such that ${\rm supp}({\left.\kern-1.2ptI_{i}\vphantom{\big{|}}\right|_{l}})={\rm supp}({\left.\kern-1.2pt\tilde{I}_{\nu(i)}\vphantom{\big{|}}\right|_{l}})$, for any $i\in\mathcal{I}_{L}$ and $l\in L$. In particular, for any two lines $l,l^{\prime}\in L$, the following diagram commutes:
[…]
    where $\nu_{l}\colon{\left.\kern-1.2ptI_{i}\vphantom{\big{|}}\right|_{l}}\in{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})\mapsto{\left.\kern-1.2pt\tilde{I}_{\nu(i)}\vphantom{\big{|}}\right|_{l}}\in{\mathcal{B}}({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big{|}}\right|_{l}})$ (and similarly for $l^{\prime}$). In other words, ${\mathbb{M}}$ and $\tilde{\mathbb{M}}_{\delta}$ have the same matched barcodes along the lines of $L$, up to interval reordering.
We now claim that multi-parameter persistence modules that are $\delta$-candidate decompositions of a given interval decomposable multi-parameter persistence module ${\mathbb{M}}$ are $\delta$ close to ${\mathbb{M}}$ in $K$, as stated in the following approximation result, which we prove in the next section.
Proposition 3.4 (Approximation result).
Let ${\mathbb{M}}$ be an interval decomposable multi-parameter persistence module. Then, any $\delta$-candidate decomposition $\tilde{\mathbb{M}}_{\delta}$ of ${\mathbb{M}}$ satisfies
${d_{\rm I}}(\tilde{\mathbb{M}}_{\delta}\cap K,{\mathbb{M}}\cap K)\leq{d_{\rm B}}(\tilde{\mathbb{M}}_{\delta}\cap K,{\mathbb{M}}\cap K)\leq\delta,$
where ${\mathbb{M}}\cap K$ stands for the module that is isomorphic to ${\mathbb{M}}$ on $K$ and the null module outside of $K$.
Note that it is possible to generalize Proposition 3.4 to modules that are not restricted to $K$ by constraining the parts of the candidate decompositions that are outside of $K$ with Kan extensions, but we stick to our formulation for the sake of simplicity.
3.2 Proof of Proposition 3.4
In this section, we prove Proposition 3.4. This section is quite technical and can thus be skipped by readers who are most interested in the general exposition. We first provide several additional definitions in Section 3.2.1, that are a bit technical yet convenient for writing down the subsequent proofs. Then, we prove a few technical results and lemmas in Section 3.2.2, and we finally prove Proposition 3.4 in Section 3.2.3.
Notations.
We first introduce a few notations: we let $(e_{1},\dots,e_{n})$ be the canonical basis of $\mathbb{R}^{n}$, and, given a set $A\subseteq\mathbb{R}^{n}$, we let ${\rm conv}(A)$ denote the convex hull of $A$. Moreover, given a hyperplane $H\subseteq\mathbb{R}^{n}$ and its two associated vectors $a_{H},b_{H}\in\mathbb{R}^{n}$ which satisfy $H=b_{H}+\{x\in\mathbb{R}^{n}:\langle x,a_{H}\rangle=0\}$, we call $a_{H}$ the codirection of $H$. When $a_{H}$ is a vector in the canonical basis of $\mathbb{R}^{n}$, i.e., there exists $i\in\llbracket 1,n\rrbracket$ such that $a_{H}=e_{i}$, we slightly abuse notation and also call $i$ the codirection of $H$.
3.2.1 Boundaries, facets and diagonal lines
In this section, we define the upper- and lower-boundaries of interval modules, as well as their so-called facets, which are convenient characterizations of the interval supports.
Definition 3.5 (Upper- and lower-boundaries).
Given an interval $I$, its upper-boundary $U[I]$ and lower-boundary $L[I]$ are defined as:
$L[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y<_{n}x\Rightarrow y\not\in I\right\},\ \ U[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y>_{n}x\Rightarrow y\not\in I\right\}$
Moreover, the boundary of ${\rm supp}(I)$ can be decomposed with $\partial{\rm supp}(I)=L[I]\cup U[I]$. See Figure 2 for an illustration.
Figure 2: Lower- and upper-boundaries of a 2-interval (Definition 3.5); and birthpoints and deathpoints $b_{x}^{I}$ and $d_{x}^{I}$ (Definition 2.10) of a point $x\in\mathbb{R}^{2}$.
When interval modules are discretely presented (see Definition 2.7), their lower- and upper-boundaries are made of flat parts, which are the faces of the corresponding rectangles forming the interval. Hence, we call facets the subsets of the lower- and upper-boundaries that are included in some hyperplanes of $\mathbb{R}^{n}$.
Definition 3.6 (Facet).
A lower (resp. upper) facet of an interval $I$ is an $(n-1)$-submanifold of $\partial{\rm supp}(I)$ written as $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap L[I]$ (resp. $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap U[I]$) for some $c\in\mathbb{R}$ and some dimension $i\in\llbracket 1,n\rrbracket$ that is called the facet codirection. In particular, the upper- and lower-boundaries of a discretely presented interval module is a locally finite union of facets.
We end this section by defining some properties of diagonal lines.
Definition 3.7 ($\delta$-regularly distributed lines filling a compact set).
@@ 3.2.1 Boundaries, facets and diagonal lines
    two diagonal lines $l,l^{\prime}\in L$ are $\delta$-consecutive (or consecutive when $\delta$ is clear) if there exists ${\overrightarrow{u}}\in\left\{0,1\right\}^{n}\backslash\left\{\mathbf{0},\mathbf{1}\right\}$ such that $l^{\prime}=l\pm\delta\cdot{\overrightarrow{u}}$.
@@
    two diagonal lines $l,l^{\prime}\in L$ are $\delta$-comparable if there exists a positive or negative vector $\overrightarrow{u}\in\mathbb{R}^{n}$ (i.e. a vector whose coordinates are all positive or all negative) with $\left\lVert\overrightarrow{u}\right\rVert_{\infty}\leq\delta$ such that $l^{\prime}=l+\overrightarrow{u}$. If $\overrightarrow{u}$ is positive (resp. negative), we write $l^{\prime}\geq l$ (resp. $l^{\prime}\leq l$).
@@
    $L$ $\delta$-fills $K$ (or fills $K$ when $\delta$ is clear) if any point of $K$ is at distance at most $\delta/2$ from some line in $L$. More formally, $K$ is included in the offset $L^{\delta/2}$.
Remark 3.8.
One can check that the $\delta$-grid of lines of a hyperrectangle $K$ (used in Definition 3.3) is $\delta$-regularly distributed and $\delta$-fills $K$.
3.2.2 Additional lemmas
In this section, we prove a few preliminary results about endpoints of interval modules, that will turn out useful for proving Proposition 3.4. The first simple, yet fundamental lemma that we will use repeatedly states that endpoints of interval modules cannot be comparable.
Lemma 3.9.
Let $l_{1},l_{2}$ be two diagonal lines and $I$ be an interval module such that the barcodes ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})$ and ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})$ are not empty. Let ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})=[b_{1},d_{1}]$ and ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})=[b_{2},d_{2}]$. Then $b_{1}$ and $b_{2}$ are not comparable, i.e., $b_{1}\not\leq_{n}b_{2}$ and $b_{1}\not\geq_{n}b_{2}$. The same is true for $d_{1}$ and $d_{2}$. In other words, the rectangles $R_{b_{1},b_{2}}$, $R_{b_{2},b_{1}}$, $R_{d_{1},d_{2}}$ and $R_{d_{2},d_{1}}$ are flat (i.e., one side has length 0).
Proof.
Figure 3: Two bars $[b_{1},d_{1}]$ and $[b_{2},d_{2}]$ of an interval module.
Stability of close barcodes.
We now show that endpoints of bars in barcodes from lines that are close should also be close. In other words, bars of the fibered barcode that are associated to lines that are close to each other must have similar length, as stated in the lemma below (see also [Lan18, Lemma 2]).
Lemma 3.10.
Let $I$ be an interval module, let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines and let $\overrightarrow{u}\in\mathbb{R}^{n}$ be a positive or negative vector (i.e., the coordinates of $\overrightarrow{u}$ are either all positive or all negative) such that $l_{2}=l_{1}+\overrightarrow{u}$. Then, the following properties hold:

### v4

@@ 2 Background
Using birthpoints and deathpoints, the $L$-fibered barcode of an interval decomposable multi-parameter persistence module $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ can be written as:
${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}=\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\right)_{l\in L}=\left(\{[b^{I_{i}}_{l},d^{I_{i}}_{l})\}_{i\in\mathcal{I}}\right)_{l\in L}.$ (6)
Remark 4.
Notice that in Remark 3, the persistence barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\{[b^{I_{i}}_{l},d^{I_{i}}_{l})\}_{i\in\mathcal{I}}$ can be seen as multisets of segments $b_{i},d_{i})$ in $\mathbb{R}^{n}\cup\left\{\boldsymbol{\infty}{}\right\}$. In particular, the diagonal line of a given segment $b_{i},d_{i})$ can be recovered from, for instance, the birthpoint $b_{i}$, and hence, without loosing any information (except for lines with trivial barcodes), we will consider the following identification:
${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}=\bigcup_{l\in L}{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\left\{{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right):l\in L\right\}.$ (7)
Geometry and stability of diagonal barcodes.
We now present two simple yet fundamental results on diagonal barcodes. The first one characterizes rectangles formed by endpoints.
Lemma 1.
Let $l_{1},l_{2}$ be two diagonal lines and $k^{I}$} be a f.p. interval module such that the barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ are not empty. Let ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)=[b_{1},d_{1})$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)=[b_{2},d_{2})$. Then, the rectangles $R_{b_{1},b_{2}}$, $R_{b_{2},b_{1}}$, $R_{d_{1},d_{2}}$ and $R_{d_{2},d_{1}}$ are flat, that is, they either have null volume, or their corners are not comparable.
Proof.
Figure 4: Two bars $b_{1},d_{1})$ and $b_{2},d_{2})$ of an interval module.
We now show that endpoints of bars in barcodes associated to lines that are close should also be close. In other words, bars of the fibered barcode that are associated to lines that are close to each other must have similar length, as stated in the lemma below; see also (43, Lemma 2).
Lemma 2.
Let $k^{I}$} be a f.p. interval module, let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines and let $\overrightarrow{u}\in\mathbb{R}^{n}$ be a positive or negative vector such that $l_{2}=l_{1}+\overrightarrow{u}$. Then, the following properties hold:
  1. (i)
    If the barcode ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)=\{[b^{I}_{l_{1}},d^{I}_{l_{1}})\}$ is not empty and satisfies $\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$, then the barcode ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ is not empty as well, and
  2. (ii)
@@
    If the barcodes ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)$ are not empty, then one has
Proof.
Item $(i)$. Since $|(d^{I}_{l_{1}})_{i}-(b^{I}_{l_{1}})_{i}|=\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$ for any index $i\in\llbracket 1,n\rrbracket$, it follows that $b^{I}_{l_{1}}\leq_{n}b^{I}_{l_{1}}+\overrightarrow{u}\leq_{n}d^{I}_{l_{1}}$. [^4] We assume here that $\overrightarrow{u}$ is positive. It should be replaced by $d_{l_{1}}^{I}-\overrightarrow{u}$ if it is negative. Thus $b^{I}_{l_{1}}+\overrightarrow{u}$ must belong to I{I} since $I$ is an interval. Hence, since $b^{I}_{l_{1}}+\overrightarrow{u}\in l_{2}$, one has ${\mathcal{B}}\left({\left.\kern-1.2pt{k^{I}}\vphantom{\big|}\right|_{l_{2}}}\right)={I}\cap l_{2}\neq\varnothing$.
Item $(ii)$. If one of the endpoints is infinite, the result holds trivially as the other endpoint has to be infinite too, so we now assume that the endpoints of the bars are all finite. Without loss of generality, assume that $l_{2}=l_{1}+\overrightarrow{u}$ where $\overrightarrow{u}$ is positive. Now, since both $d_{l_{2}}^{I}$ and $d_{l_{1}}^{I}+\overrightarrow{u}$ belong to $l_{2}$, they are comparable, so one has either $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}$ or $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. However, the first possibility would lead to $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}>_{n}d_{l_{1}}^{I}$, hence $d_{l_{1}}^{I}$ and $d_{l_{2}}^{I}$ would be (strictly) comparable in $\mathbb{R}^{n}$, which contradicts Lemma 1. Thus, one must have $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. Furthermore, and using the exact same arguments, $d^{I}_{l_{2}}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}$ is on $l_{1}$, and one must have $d_{l_{2}}^{I}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}\geq_{n}d_{l_{1}}^{I}$. Finally, by combining the two previous inequalities, one has:
3 Computing candidate decompositions with the MMA algorithm
@@
In this section, we present our family of descriptors for multi-parameter persistence modules, defined as candidate decompositions into interval summands and we identify a specific subfamily that we call approximate decompositions (Definition 19), in Section 3.1. Then, we show how practical computations of instances of such candidate decompositions can be done with our MMA algorithm in Sections 3.2 and 3.3.
3.1 Candidate and approximate decompositions
@@
Our candidate decompositions depend on $\delta$-grids of lines, that we now define.
Definition 16 ($\delta$-grid of lines).
@@
where $l_{\delta\cdot u}:=\delta\cdot u+\boldsymbol{e}_{\Delta}{\mathbb{R}}$ is the diagonal line with direction vector $\boldsymbol{e}_{\Delta}=[1,\dots,1]^{T}\in\mathbb{R}^{n}$ passing through $\delta\cdot u$.
Several new definitions can be introduced from grids of lines, which will turn useful either in the definition of our MMA algorithm, or in the corresponding theoretical proofs.
Definition 17 ($\delta$-regularly distributed lines filling a compact set).
@@
[…]
@@
    Two diagonal lines $l,l^{\prime}\in L$ are $\delta$-comparable if there exists a positive or negative vector $\overrightarrow{u}\in\mathbb{R}^{n}$ with $\left\lVert\overrightarrow{u}\right\rVert_{\infty}\leq\delta$ such that $l^{\prime}=l+\overrightarrow{u}$. If $\overrightarrow{u}$ is positive (resp. negative), we write $l^{\prime}\geq l$ (resp. $l^{\prime}\leq l$).
@@
    $L$ $\delta$-fills $K$ (or fills $K$ when $\delta$ is clear) if any point of $K$ is at distance at most $\delta/2$ from some line in $L$. In other words, $K$ is included in the offset $L^{\delta/2}$.
Our candidate decompositions of a given multi-parameter persistence module $\mathbb{M}$ are, roughly speaking, interval decomposable modules with fibered barcodes containing the one of $\mathbb{M}$ on a $\delta$-grid of lines. Before going into the definition of our estimator candidate, we introduce a compactness assumption, that directly follows Remark 1.
Definition 18 (Module compactness).
We say that a rectangle $K=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]\subseteq\mathbb{R}^{n}$} compactly characterizes an $n$-parameter persistence module $\mathbb{M}$ if restricting $\mathbb{M}$ to $K$ preserve information, or, more formally, if $\mathrm{Lan}_{n}\,{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}}\simeq\mathbb{M}$.
Definition 19 (Candidate and approximate decompositions).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module. Let $K$ be a compact set that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}=\{x\in\mathbb{R}^{n}:d_{\infty}(x,K)\leq\delta\}$, where $d_{\infty}$ stands for the $\left\lVert\cdot\right\rVert_{\infty}$ distance. A multi-parameter persistence module $\tilde{\mathbb{M}}_{\delta}$ is called a $\delta$-candidate decomposition of $\mathbb{M}$ if: